OpenStudy (anonymous):

Solve for x. 4 • 7 – 4 • 10 = 4(x – 10) A. –13 B. –6 C. 4 D. 7

OpenStudy (anonymous):

28-40=4x-40 -12=4x-40 28=4x 7=x